EXPERIMENTAL APPROACHES TO THE MEASUREMENT OF BROAD RESONANCES IN THE MI-CROWAVE, MILLIMETER, AND SUBMILLIMETER SPECTRAL REGIONS

\underline{FRANK C. DE LUCIA} and ANDREY MESHKOV, Department of Physics, The Ohio State University, Columbus OH 43210.

Microwave spectroscopists have taken advantage of the narrowness of resonances in the gas phase to separate spectral resonances from broader variations in the system power. However, there are a growing number of physical phenomena, especially in the condensed phase, which have either broad resonances or ’continua’ associated with them. In this talk approaches to measurements of these phenomena will be discussed and examples provided.